Sol Summary Report – Mar 31st

Crew 208 Sol Summary Report 31-MAR-2019

Sol: 3

Summary Title: Welcome to MDRS, NAU!

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Crew 208 continued to 3D print our drone map.

Crew 208 completed MDRS orientation, rover and spacesuit training today.

Our Crew 208 team photo was a group selfie taken via drone (i.e., a “dronie”).

GHO used Crew 208’s Insta-Pot to prepare a vegetarian lunch meal.

COM gave a tour of MDRS to NAU graduate students from 1400-1600.

XO + ENG + GHO + HSO drove rovers towards Candor Chasma for drone flights today.

XO + David traveled to Hanksville to refill the Greenhab water tank.

Look Ahead Plan: Continue 3D printing maps of MDRS in Hab. Start data collection for thoracentesis study. Begin sim this PM. David to deliver Crew’s 208 food supply on Sol 4.

Anomalies in work: Internet was offline all-day until 2040; Opportunity rover experienced battery overheating on Cactus Road 1104 requiring ATV towing by David.

Weather: Sunny

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab, HSO Safety Checklist, Journalist, EVA Request

Support Requested: Please repair Opportunity rover

Journalist Report – Mar 31st

Journalist Report

Date: March 31, 2019

Author: Erika Rydberg

Crew 208 Reporting for Duty on Mars

Today we experienced our first full day as a crew, we were very excited to have five people on board for this journey and experience – Amanda, David, Julielynn, Dean and myself, Erika. We began the day with a shared breakfast of cereal, dried fruit, our preferred caffeinated beverages, and of course water, as one must never forget plenty of water in this dry environment. We surely expend more energy than we even realize on a daily basis with the work and exploring the vast landscapes around MDRS, and water is key to sustaining our activity and our well-being. Our community building has begun as we’ve started to share cooking duties and the adventures of cooking with the various unique supplies of our home. You can do a lot with a little kitchen expertise: we’ve had soup, pasta with pork and tomato sauce and beans and rice today, we’re lucky to have curious cooks in our crew. Over our delicious food today conversation was bountiful and we started the day with a fantastic conversation about space, becoming an astronaut, and the future ahead for some of the new companies creating opportunities to visit space.

One of our 3D printers continued progressing smoothly making a 3D map of MDRS first thing this morning. We’re happy to have some of our machinery working well in this complex dry environment.

For the three new members to the Hab at 10:00 sharp we reported to be trained on driving our rovers and the process of EVAs and we went on a small trip to experience more of the vast and curious landscapes we find ourselves in. The geology and landscape is incredible, one moment you’re looking at mounds of varying red, green and black land and then in the next moment, we’ve found a deep vast canyon. Upon returning from we took a quick crew “dronie” (a selfie with a drone), and then we were able to try on and use our space suits for the first time and attempt to see what it is like to maneuver first of all our bodies and then second of all our rovers while existing within a complex suit. On one hand, the suit feels cumbersome, but on another, the suit provides comfort through both safety and protection.

Julielynn greeted graduate students from Northern Arizona University and brought them through a tour of the Hab and MDRS. The rest of Crew 208 meanwhile, ran through some of our duties. We then adventured out again in our rovers for one last trip of fresh air and some more drone flying. One of our rovers got a bit stuck on the way back, but with the help of a tow, we were able to return safe and sound to our home, the Hab ready to see what lies ahead.

Operations Report – March 30th

Crew 208 Operations Report 30-MAR-2019

SOL: 2

Name of person filing report: Amanda Manget

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 15

Solar – SOC 99% @ 1900 (before David turned on the generator for the night)

Diesel Reading – 48%

Propane Reading – 56%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 228 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 114 gallons

Static to Loft Pump used – Yes

Water Meter: 142228.4 gallons @ 2000

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 85.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 54.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Not Used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) N/A

Reason for use: Not Used

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where?

Hab car was driven to Hanksville to pick up water for Crew 208

Har car was driven to Grand Junction, CO to pick up 2 members of Crew 208

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read >12.5 volts

Summary of Hab operations: Checked SOC of the electrical system as follows:

0930 SOC 100%, generator turned off by David
1500 SOC 100%, generator off
1900 SOC 99%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: None (Note: observatory is closed)

Summary of health and safety issues: None

Greenhab Report – March 30th

Crew 208 GreenHab Report – 30-March-2019

GreenHab Officer: Erika Rydberg

Environmental control:

Ambient with the door open

Shade cloth (80%) on

Average temperatures: 22 degrees Celsius

Low: 12
High: 29

Hours of supplemental light: 0

Daily water usage for crops: 6 gallons

Water in Blue Tank — 114 gallons

Time(s) of watering for crops:10:45 and 16:55

Changes to crops: None

Narrative: First day for me in the Greenhab, I garden at home, so it was nice to see some good growth coming from vegetables in this environment. Herbs are abundant.

Harvest: 2 grams basil

Support/supplies needed: None

Sol Summary – March 30th

Crew 208 Sol Summary Report 30-MAR-2019

Sol: 2

Summary Title: Welcome to MDRS, Crew 208!

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

GHO arrived yesterday PM.

COM trained GHO on filling out 3D print logs.

Crew 208 completed 3D printing in the Hab today for the thoracentesis training study.

XO continued to process MDRS maps from drone images. GHO prepared 3D print files of our drone maps and started printing our 1st 3D map of MDRS. XO + GHO conducted several drone flights today.

XO + David traveled to Hanksville to refill our water resources at MDRS.

COM trained incoming GHO on GreenHab duties and reports. XO trained incoming ENG on reporting duties.

David drove CrewCar to GJT to meet Crew 208 ENG + HSO + transport to MDRS.

Crew 208 will drive into Hanksville tonight for dinner after reports have been submitted to CAPCOM.

Look Ahead Plan: Continue 3D printing maps of MDRS in Hab. MDRS orientation and training to commence on Sol 2 and finish on Sol 3. COM to give a tour to PS students from 1400-1600 on Sol 3. Begin sim in PM of Sol 3.

Anomalies in work: None

Weather: Sunny

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab

Support Requested: None

Operations Report – March 29th

Crew 208 Operations Report 29-MAR-2019

SOL: 1

Name of person filing report: Dean Jin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 13

Solar – SOC 88% @ 1930 (before David turned on generator for the night)

Diesel Reading – 48%

Propane Reading – 58%

Ethanol Free Gasoline – 2 gallons

Water (auxiliary tank) – 0 gallons (Note: Not in Use)

Water (static tank) – 124 gallons

Auxiliary to Static tank transfer – No

Gallons transferred: 0

Water in GreenHab – 120 gallons

Static to Loft Pump used – Yes

Water Meter: 142212.8 gallons @ 2000

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours: 85.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: No

Hours: 54.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Not Used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) N/A

Reason for use: Not Used

Oil Added? No

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: None

HabCar used and why, where? Not Used

CrewCar used and why, where? Not Used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EV suits normal, all batteries read > 12.5 volts

Summary of Hab operations: Checked SOC of electrical system as follows:

0800 SOC 100%, generator turned off by David

1100 SOC 100%, generator off

1500 SOC 96%, generator off

1800 SOC 90%, generator off

1930 SOC 88%, generator turned on by David

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAMM operations: Nothing to report

Summary of any observatory issues: None (Note: observatory is closed)

Summary of health and safety issues: None

Cheers,

Julielynn Wong, MD, MPH, FACPM

GreenHab Report – March 29th

Crew 208 GreenHab Report – 29-MAR-2019

GreenHab Officer: Julielynn Wong

Environmental control:

Ambient (no heating or/cooling)

Shade cloth (80%) on

Average temperatures: 25 degrees Celsius

Low 21 degrees Celsius

High 32 degrees Celsius

Hours of supplemental light: None

Daily water usage for crops: 4 gallons

Water in Blue Tank— 120 gallons

Time(s) of watering for crops: 1100 1515

Changes to crops: None

Narrative: Awaiting arrival of Crew 208 GHO

Harvest: None

Support/supplies needed: None

Cheers,

Julielynn Wong, MD, MPH, FACPM

Sol Summary – March 29th

[category sol-summary

Crew 208 Sol Summary Report 29-MAR-2019

Sol: 1

Summary Title: Welcoming Crew 208 GHO to MDRS!

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

COM completed the end of mission food survey for Crew 207 and submitted the Crew 207 Mission Summary to Mission Support.

Crew 208 used one 3D printer in the Hab to prepare for our thoracentesis training study.

XO continues to fulfill ENG duties. COM continues to fulfill GHO duties.

XO travelled to GJT today to bring GHO to MDRS this PM.

Look Ahead Plan: David to drive CrewCar to GJT on Sol 2 to meet Crew 208 ENG + HSO for transport to Bull Mountain Market + MDRS. MDRS orientation and training to commence on Sol 2 and finish on Sol 3. Begin sim in PM of Sol 3.

Anomalies in work: None

Weather: Sunny with high winds

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab

Support Requested: None

Cheers,

Julielynn Wong, MD, MPH, FACPM